10. Considering a recession may happen next year. Commercial banks are more conservative in their lending policies and start7. Refer to the table below. The required reserve ratio is 25%. If the First Charter Bank is meeting its reserve requirement6. First National Bank has assets of $500,000 and liabilities of $400,000. First National Banks equality (net capital) is A.
Answer #1

10.

D

When banks start holding excess reserve, then less loan and lesser deposit creation will take place. As a result, money supply will be lower.

---------------

7.

B

Working note:

Loans = 1200 - required reserve = 1200 - 25%*800

Loans = $1000

------------------

8.

A

Working note:

Required reserve ratio = (400000 - 100000)/2000000

Required reserve ratio = 15%

------------------

6.

C

Working note:

Net capital = Assets - liability = 500000 - 400000

Net Capital = $100000